1. What is the boiling point and freezing point of a solution containing 6 g of



glycerol (C3H8O3) in 25 g of water?



2. What is the molecular mass of a solute in a solution containing 3.8 g of



nonelectrolyte in 420 g of water which freezes at -23 oC?

1
Expert's answer
2022-04-20T10:26:04-0400

ΔT = Kb*molality

Moles = 6 / 92 = 0.065 mol

Molality = 0.065 / 0.025 = 2.6

ΔT = 0.51 * 2.6 = 1.33

T2 = 100 + 1.33 = 101.330C

 

ΔT = Kf*molality

ΔT = 1.86 * 2.6 = 4.84

T2 = 0 – 4.84 = -4.84